To construct a 99% confidence interval where o is known, the correct critical value is 1.96.

True or false
Answer #1

False

As the sigma is known we use z test The critical value at 99% is 2.58

Given CI level is 99%, hence α = 1 - 0.99 = 0.01
α/2 = 0.01/2 = 0.005, Zc = Z(α/2) = 2.58